OpenStudy (anonymous):

How do i solve this? x^3+x^2-4x-4=0

jimthompson5910 (jim_thompson5910):

x^3+x^2-4x-4=0 (x^3+x^2)+(-4x-4)=0 x^2(x+1)+(-4x-4)=0 x^2(x+1)-4(x+1)=0 (x^2-4)(x+1)=0 (x-2)(x+2)(x+1)=0 x-2=0, x+2=0, or x+1=0 x=2, x=-2, or x=-1

OpenStudy (anonymous):

On the third step, why do you put x^2(x+1) ?

jimthompson5910 (jim_thompson5910):

I factored out x^2 from x^3+x^2

jimthompson5910 (jim_thompson5910):

I did this so I could get two copies of x+1 (afterwards I factored -4 from -4x-4 to get another copy of x+1)

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

Not really. Solving equations, has always been the hardest for me.

jimthompson5910 (jim_thompson5910):

well going from x^3+x^2-4x-4=0 to (x-2)(x+2)(x+1)=0, I'm factoring the left side once everything is factored, I then use the zero product property to solve for x.

OpenStudy (anonymous):

Cant you use the distance formula?

jimthompson5910 (jim_thompson5910):

the distance formula won't work too well in this setting, so I'm not sure why you're thinking of that (unless this is for another problem?)